OpenStudy (anonymous):

Solve the following equation: -2 (x - 3) = -8

OpenStudy (anonymous):

start by dividing both sides of equation by -2

OpenStudy (anonymous):

No lol.... first you distribute the -2 by multiplying -2*x and -2*-3

OpenStudy (anonymous):

x=7 right ?

OpenStudy (anonymous):

Yesmam (:

OpenStudy (anonymous):

divide both sides by -2 gives: x - 3 = 4 => x = 4 + 3 = 7